OpenStudy (rywisely):

Simplifying radicals: 6√500x^3

OpenStudy (zzr0ck3r):

\(6\sqrt{500x^3}\) ?

OpenStudy (zzr0ck3r):

yes no?

OpenStudy (rywisely):

Yes, that's right.

OpenStudy (zzr0ck3r):

\(6\sqrt{500x^3}=6\sqrt{25*4*5*x^3}=6\sqrt{25}\sqrt{4}\sqrt{5}\sqrt{x^3}\\=6*5*2*\sqrt{5}x^{\frac{3}{2}}=60\sqrt{5}x^{\frac{3}{2}}\)

OpenStudy (zzr0ck3r):

Let me know if any step confuses you.

OpenStudy (rywisely):

Thank you so much! I'll let you know if I'm confused.

OpenStudy (rywisely):

How did you get the part that says, "6*5*2*\[\sqrt{5}x\frac{ 3 }{ 2 }\]?

OpenStudy (anonymous):

\[60 x \sqrt{5 x} \]

OpenStudy (rywisely):

I'm confused on how you get those numbers to multiply.